OpenStudy (anonymous):

how do i simplify 3 cos^2 x + 3 cos^2 x tan^2 x

OpenStudy (anonymous):

\[\tan^2x=\frac{\sin^2x}{\cos^2x}\] so \[\cos^2x\tan^2x=\sin^2x\] \[3\cos^2x+3\sin^2x=3\left(\cos^2x+\sin^2x\right)=\cdots\]

OpenStudy (kaederfds):

and remember \[\cos^{2}x + \sin^{2}x = 1\] :)

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

3